# DiaHalu: A Dialogue-level Hallucination Evaluation Benchmark for Large Language Models

### v2

@@ Abstract
Since large language models (LLMs) achieve significant success in recent years, the hallucination issue remains a challenge, and numerous benchmarks are proposed for hallucination detection. Nevertheless, some of these benchmarks are not naturally generated by LLMs but are intentionally induced. Also, many merely focus on the factuality hallucination while ignoring the faithfulness hallucination. Additionally, although dialogue pattern is more widely utilized in the era of LLMs, current benchmarks only concentrate on sentence-level and passage-level hallucination. In this study, we propose DiaHalu, the first dedicated dialogue-level hallucination evaluation benchmark for LLMs to our knowledge. Initially, we integrate the collected topics into system prompts and facilitate a dialogue between two LLMs. Subsequently, we manually modify the contents that do not adhere to human language conventions and then have LLMs re-generate, simulating authentic human-machine interaction scenarios. Finally, professional scholars annotate all the samples in the dataset. DiaHalu covers four common multi-turn dialogue domains and five hallucination subtypes, extended from factuality and faithfulness hallucination. Experiments through some well-known LLMs and detection methods on the dataset show that DiaHalu is a challenging benchmark, holding significant value for further research [^1] https://github.com/141forever/DiaHalu.
@@ 1 Introduction
Despite many advantages of large language models, the issue of hallucination remains a primary concern (Ji et al., 2023; Zhang et al., 2023c). Hallucination predominantly centers on the inclination of LLMs to generate nonsensical or untruthful contents for specific sources (Wang et al., 2023a). Therefore, the occurrence of hallucination poses risks to the application of large language models in various real-world scenarios (Omiye et al., 2023; Wu et al., 2023b; Chen et al., 2024).
Given the aforementioned risks, hallucination detection emerges as a highly crucial task. In recent years, researchers propose numerous benchmarks for hallucination detection task (Li et al., 2023b; Guan et al., 2023; Manakul et al., 2023; Yang et al., 2023). Nevertheless, several problems persist in these benchmarks. (1) Not naturally generated. One pitfall of existing benchmarks for detecting LLMs’ hallucination is that the hallucinated contents are typically induced via manually designed trigger prompts (Li et al., 2023b), while not naturally generated by LLMs as in daily usage (Liu et al., 2022). (2) Merely focusing on factuality hallucination. Most benchmarks merely focus on detecting factuality hallucination (Guan et al., 2023), with few datasets that can demonstrate faithfulness hallucination (Huang et al., 2023a) (Figure 1). (3) Only concentrating on sentence-level and passage-level. Researchers propose many sentence-level (Manakul et al., 2023; Zhao et al., 2023c) and passage-level (Yang et al., 2023; Feng et al., 2023b) hallucination detection benchmarks. However, the dialogue pattern has broader and more widespread applications in LLMs. Although it is rarely mentioned in previous researches, dialogue-level hallucination detection is equally essential.
@@ 2.2 Hallucination Detection
Current methods for hallucination detection (Tonmoy et al., 2024) can mainly be divided into four categories. (1) Model-Based. This method involves having the language models perform a classification task to determine whether hallucinated contents are present (Zhao et al., 2023a; Maharaj et al., 2023). (2) Retrieval-Based. For the limited knowledge within the parameters of language models, we can detect hallucination by extracting or retrieving relevant knowledge from external knowledge graphs (Martino et al., 2023; Chen et al., 2023c) or web information sources (Béchard and Ayala, 2024). (3) Sample-Based. Another feasible method is to rewrite the generated contents to evaluate the consistency (Manakul et al., 2023; Zhao et al., 2023c; Zhang et al., 2023a). (4) Proxy-Based. The mainstream view of this method (Zhang et al., 2023b; Gupta et al., 2024) is that ‘the lower the probability of generating a token, the more likely a model is to produce hallucination’.
@@ 4.3 Human Annotation
Table 5: The results of CoT and retrieval techniques on the three closed-source LLMs. The $\uparrow$ and $\downarrow$ indicate whether CoT and retrieval can promote improvements in F1 score. We also provide the detection results of FaithCritic.
@@ 5.2 Metrics
For hallucination detection, we use standard binary classification to determine whether there exists hallucination (Table 4). Meanwhile, we also conduct more fine-grained hallucination-type recognition to judge the specific subtype of hallucination and use micro-F1 score for all hallucination categories (Table 6). Appendix A.8 (II. Metrics Calculation) provides more thorough explanations.
@@ 5.3 Main Results
Table 6: Fine-grained hallucination-type recognition F1 scores for three LLMs. ‘NF’, ‘Ic’, ‘Ir’, ‘Ov’ and ‘RE’ stand for Non-factual, Incoherence, Irrelevance, Overreliance and Reasoning Error, respectively. ‘ALL’ represents micro-f1 of all hallucination subtypes.
@@
5.4 Chain-of-Thought and Retrieval for Detection
Chain-of-Thought (CoT) and Retrieval are two important techniques for enhancing the ability of LLMs. In this section, we test whether these two techniques can improve the effectiveness of hallucination detection in Table 5. More details are depicted in Appendix A.11.
The experimental results indicate that both methods have facilitating effects on hallucination detection. However, Gemini1.5 PRO and ChatGPT3.5 with CoT show a decrease of around 1.00 F1 points in the chit-chat domain. We believe that these two models inherently lack the ability to recognize faithfulness hallucination, and additional CoT contents introduce noise to their judgments.
@@ 5.5 Fine-grained Hallucination-type Recognition
Table 6 shows fine-grained hallucination-type recognition results for three open-source LLMs. We can conclude that ChatGPT3.5 fails the recognition of almost all labels. To some extent, Gemini1.5 PRO and GPT4 have the ability to recognize factuality hallucination and reasoning errors, but they have lower F1 scores for the three subtypes of faithfulness hallucination. This result reveals that faithfulness hallucination remains a pressing issue for LLMs.
@@ Acknowledgements
Due to the constrictions of the double-blind review mechanism, we will acknowledge all contributors and institutions as soon as our paper is accepted.
@@ A.3 The Distribution of the Sources for Dialogue Topics
The number and proportion of all dialogue topics in DiaHalu across the 10 topic sources are shown in Figure 6.
@@ A.12 Future Works
Here is a sentence-level hallucination example from the dataset FactCHD (Chen et al., 2023c). ‘User: Can you tell me which mountain range is longer, the Alps or the Pyrenees? LLMs: The Pyrenees are longer than the Alps.’ In this sentence-level example, the LLM only responds to the user’s question with one sentence. We need to assess whether there are hallucination in the single sentence generated by the LLM.
@@
is the phenomenon that LLMs tend to accumulate hallucination rather than self-correcting during the generation process (Zhang et al., 2023c). Some previous works validate this phenomenon (Azaria and Mitchell, 2023; Ang et al., 2023). In our benchamark, there is a noticeable issue of hallucination snowballing. Through the experimental results, we also display such phenomenon. Eliminating hallucination snowballing in LLMs is extremely urgent in the future.

### v3

@@ Abstract
Though large language models (LLMs) achieve significant success in recent years, the hallucination issue remains a challenge, and numerous benchmarks are proposed for hallucination detection. Nevertheless, some of these benchmarks are not naturally generated by LLMs but are intentionally induced. Also, many merely focus on the factuality hallucination while ignoring the faithfulness hallucination. Additionally, although dialogue pattern is more widely utilized in the era of LLMs, current benchmarks only concentrate on sentence-level and passage-level hallucination. In this study, we propose DiaHalu, the first dedicated dialogue-level hallucination evaluation benchmark for LLMs to our knowledge. Initially, we integrate the collected topics into system prompts and facilitate a dialogue between two LLMs. Subsequently, we manually modify the contents that do not adhere to human language conventions and then have LLMs re-generate, simulating authentic human-machine interaction scenarios. Finally, professional scholars annotate all the samples in the dataset. DiaHalu covers four common multi-turn dialogue domains and five hallucination subtypes, extended from factuality and faithfulness hallucination. Experiments with the well-known LLMs and detection methods show that DiaHalu is a challenging benchmark, holding significant values for further research [^1] https://github.com/ECNU-ICALK/DiaHalu.
@@ 1 Introduction
Despite many advantages of LLMs, the issue of hallucination remains a primary concern (Ji et al., 2023; Zhang et al., 2023c). Hallucination predominantly centers on the inclination of LLMs to generate nonsensical or untruthful contents for specific sources (Wang et al., 2023a). Therefore, the occurrence of hallucination poses risks to the application of LLMs in various real-world scenarios (Omiye et al., 2023; Wu et al., 2023b; Chen et al., 2024).
Given the aforementioned risks, hallucination detection emerges as a highly crucial task. In recent years, researchers propose numerous benchmarks for hallucination detection task (Li et al., 2023b; Guan et al., 2023; Manakul et al., 2023; Yang et al., 2023). Nevertheless, several problems persist in these benchmarks. (1) Not naturally generated. One pitfall of existing benchmarks for detecting LLMs’ hallucination is that the hallucinated contents are typically induced via manually designed trigger prompts (Li et al., 2023b), while not naturally generated by LLMs as in daily usage (Liu et al., 2022). (2) Merely focusing on factuality hallucination. Most benchmarks merely focus on detecting factuality hallucination (Guan et al., 2023), with few datasets that can demonstrate faithfulness hallucination (Huang et al., 2023a) (Figure 1). Factuality hallucination emphasizes the discrepancy between generated contents and real-world facts. Faithfulness hallucination refers to the divergence of generated contents from user instructions or other inputs, as well as self-consistency within the generated contents, which tends to be ignored. (3) Only concentrating on sentence-level and passage-level. Researchers propose many sentence-level (Manakul et al., 2023; Zhao et al., 2023c) and passage-level (Yang et al., 2023; Feng et al., 2023b) hallucination detection benchmarks. However, the dialogue pattern has broader and more widespread applications in LLMs. More unique types of hallucination in dialogues make them more challenging to be detected (detailed explanations in Appendix A.12). Although it is rarely mentioned in previous researches, dialogue-level hallucination detection is equally essential.
@@ 2.2 Hallucination Detection
Current methods for hallucination detection (Tonmoy et al., 2024) can mainly be divided into four categories. (1) Model-based. This method involves having the language models perform a classification task to determine whether hallucinated contents are present (Zhao et al., 2023a; Maharaj et al., 2023). (2) Retrieval-based. For the limited knowledge within the parameters of language models, we can detect hallucination by extracting or retrieving relevant knowledge from external knowledge graphs (Martino et al., 2023; Chen et al., 2023c) or web information sources (Béchard and Ayala, 2024). (3) Sampling-based. Another feasible method is to rewrite the generated contents to evaluate the consistency (Manakul et al., 2023; Zhao et al., 2023c; Zhang et al., 2023a). (4) Uncertainty-based. The mainstream view of this method (Zhang et al., 2023b; Gupta et al., 2024) is that ‘the lower the probability of generating a token, the more likely a model is to produce hallucination’.
@@ 4.3 Human Annotation
w/ one-shot 80.43 21.51 33.94$\uparrow$ 60.87 37.33 46.28$\uparrow$ 70.91 39.39 50.65$\uparrow$ 73.91 52.31 61.26$\uparrow$ 71.97 36.13 48.11$\uparrow$
@@
w/ one-shot 40.00 1.16 2.26$\uparrow$ 42.85 4.00 7.31$\uparrow$ 46.15 6.06 10.71$\uparrow$ 56.25 6.92 12.32$\uparrow$ 48.78 4.20 7.73$\uparrow$
@@
w/ one-shot 81.42 33.14 47.11$\uparrow$ 71.87 30.67 42.99$\downarrow$ 72.22 26.26 38.52$\uparrow$ 73.11 66.92 69.87$\uparrow$ 75.09 40.55 52.66$\uparrow$
@@
Table 5: The results of CoT, the one-shot settings and the retrieval technique on the three closed-source LLMs. The $\uparrow$ and $\downarrow$ indicate whether these can promote improvements in F1 score. We also provide the detection results of FaithCritic with the retrieval technique.
@@ 5.2 Metrics
For hallucination detection, we use standard binary classification to determine whether there exists hallucination (Table 4). We utilize binary classification evaluation metrics: Precission, Recall and F1. The positive label for this classification task is set as "Halu". Meanwhile, we also conduct more fine-grained hallucination-type recognition to judge the specific subtype of hallucination and use micro-F1 score for all hallucination categories (Table 6). Appendix A.8 (II. Metrics Calculation) provides more thorough explanations.
@@ 5.3 Main Results
w/ one-shot 31.34 21.95 5.41 8.33 60.00 34.11
Table 6: Fine-grained hallucination-type recognition F1 scores for three LLMs. ‘NF’, ‘Ic’, ‘Ir’, ‘Ov’ and ‘RE’ stand for Non-factual, Incoherence, Irrelevance, Overreliance and Reasoning Error, respectively. ‘ALL’ represents micro-f1 of all hallucination subtypes. We also incorporate a sample in the prompt to assist with the judgment. If the results in the one-shot setting show improvement, we bold them.
@@
5.4 Chain-of-Thought, One-shot and Retrieval for Detection
Chain-of-Thought (CoT) and Retrieval are two important techniques for enhancing the ability of LLMs. In this section, we test whether these two techniques can improve the effectiveness of hallucination detection in Table 5. More details are depicted in Appendix A.11. Meanwhile, we also incorporate a sample based into the vanilla prompt to test the few-shot capabilities of these three language models.
The experimental results indicate that all these three methods have facilitating effects on hallucination detection. However, Gemini1.5 PRO and ChatGPT3.5 with CoT show a decrease of around 1.00 F1 points in the chit-chat domain. We believe that these two models inherently lack the ability to recognize faithfulness hallucination, and additional CoT contents introduce noise to their judgments.
@@ 5.5 Fine-grained Hallucination-type Recognition
Table 6 shows fine-grained hallucination-type recognition results for three open-source LLMs. We can conclude that ChatGPT3.5 fails the recognition of almost all labels. To some extent, Gemini1.5 PRO and GPT4 have the ability to recognize factuality hallucination and reasoning errors, but they have lower F1 scores for the three subtypes of faithfulness hallucination. This result reveals that faithfulness hallucination remains a pressing issue for LLMs. In the one-shot setting, all three models show improved accuracy in recognizing most labels. However, this setting also introduces some sample noise that interferes with the models’ judgments, such as GPT4 decreases its recognition F1 of the ‘Ic’ label.
@@ Acknowledgements
This research is funded by the National Science and Technology Major Project (No. 2021ZD0114002), the National Nature Science Foundation of China (No. 62477010, No. 62307028), the Science and Technology Commission of Shanghai Municipality Grant (No. 22511105901, No. 21511100402), and the Shanghai Science and Technology Innovation Action Plan (No. 23ZR1441800 and No. 23YF1426100).
We would like to thank all the reviewers for providing suggestions to improve our work.
@@ A.3 The Distribution of the Sources for Dialogue Topics
The number and proportion of all dialogue topics in DiaHalu across the 10 topic sources are shown in Figure 6. We use the hot topics of the social media, such as Facebook and Twitter, to obtain more topics in data generation process. GPT4 is used to augment and elaborate on the topics, scenes, and user behaviors. Take MultiWOZ dataset (which covers 7 real-life scenarios) as an example, we generate more real-life task-oriented dialogue scenarios via GPT4. As for the DSTC dataset with a focus on bus routes, we apply GPT4 to generate more potential user behaviors related to buses.
@@ A.5 Generation Details and Data Format
Humans usually respond with flexible and brief expressions in dialogues, so we remove some GPT-specific generation patterns to better simulate natural human language. These patterns include special phrases like "I am an AI, so I cannot answer the corresponding question", detailed sectional explanations of a concept or method, excessively long sentences, and so on. Chit-Chat and Reasoning dialogue can also represent human-computer interaction. According to our preliminary studies, the formatting errors and GPT-specific generation patterns (mentioned above) are less likely to occur in these two types of dialogues, thus we skip the manual examination of whether A’s response adheres to human language, and focus on other examinations such as topic diversity.
@@ A.12 Future Works
Here is a sentence-level hallucination example from the dataset FactCHD (Chen et al., 2023c). ‘User: Can you tell me which mountain range is longer, the Alps or the Pyrenees? LLMs: The Pyrenees are longer than the Alps.’ In this sentence-level example, the LLM only responds to the user’s question with one sentence. We need to assess whether there are hallucination in the single sentence generated by the LLMs.
@@
is the phenomenon that LLMs tend to accumulate hallucination rather than self-correcting during the generation process (Zhang et al., 2023c). Some previous works validate this phenomenon (Azaria and Mitchell, 2023; Ang et al., 2023). In our benchamark, there is a noticeable issue of hallucination snowballing. Through the experimental results, we also display such phenomenon. This is because that LLMs are unable to perform timely self-check during generation, leading to the accumulation of hallucinations in multi-turn dialogues. Eliminating hallucination snowballing in LLMs is extremely urgent in the future.